anel4617:

Sunfresh Bakery makes Italian bread that costs​ $1.34 per loaf. Past experience shows that 8 percent of the loaves will spoil and have to be discarded. Assuming Sunfresh wants a 45 percent markup based on cost and produces 250​ loaves, what should each loaf of bread sell​ for?

Extrinix:

so to solve this, take the cost ($1.34) and increase it by 45%, then take the amount of loaves of bread (250) and divide by the answer you got for 1.34+45% [of 1.34]

Extrinix:

\(\text{so do 1.34} \times \text{0.45 and then do} \\ \dfrac{250}{1.34 \times 0.45}\)

Extrinix:

or just do \(\dfrac{250}{1.34 \times 0.45}\)